OpenStudy (dtan5457):

How do I find the asymptotes of -3/x and 11/x-9+9? without graphing?

OpenStudy (solomonzelman):

wait you have, \(\large\color{black}{ y=\frac{\LARGE -3 }{\LARGE x} }\) and \(\large\color{black}{ y=\frac{\LARGE 11 }{\LARGE x-9} +9 }\) correct?

OpenStudy (dtan5457):

yes

OpenStudy (solomonzelman):

the asymptote will be the value that x can't equal. What value can't x equal, when you have: \(\large\color{black}{ y=\frac{\LARGE -3 }{\LARGE x} }\) ?

OpenStudy (solomonzelman):

( number/0 is undefined)

OpenStudy (dtan5457):

0

OpenStudy (solomonzelman):

yes.

OpenStudy (dtan5457):

so x and y asymptotes are 0?

OpenStudy (solomonzelman):

in this case yes.

OpenStudy (dtan5457):

to compare -3/x and 1/x, would it just be that -3/x is increasing while 1/x is decreasing?

OpenStudy (solomonzelman):

I am still typing, I dissapeared for some reason

OpenStudy (dtan5457):

i know, the site is messed up lately

OpenStudy (solomonzelman):

\(\large\color{blue}{ y=\frac{\LARGE -3 }{\LARGE x} }\) x=0 is the vertical asymptote, b/c: it y will be completely different as x approaches 0 from left and right side (of the y axis) when you plug in very small negative numbers you get a (-3/-0.001) which becomes (3/0.001) and that is 3000. NOW LETS, approach zero (from left side), a bit closer, (-3/-0.00001) >>> 3/0.00001 >>> 3000000 so as x goes closer and close to zero from the left (of the y axis), the y is infinity.

OpenStudy (solomonzelman):



OpenStudy (solomonzelman):

So, the opposite is for x approaching zero from the right side. \(\large\color{blue}{ y=\frac{\LARGE -3 }{\LARGE x} }\) (-3 / 0.01) >>> -300 (-3 / 0.0001) >>> -30,000 (-3 / 0.000001) >>> -3,000,000

OpenStudy (solomonzelman):

Same thing, but the other direction, as x gets closer and closer to zero from the right of the x-axis, the y value becomes a very large negative i.e. \(\large\color{blue}{-\infty }\)

OpenStudy (solomonzelman):